[Brahms-l] Shift report 20040428 16:00-24:00

From: Apache <apache@bnl.gov>
Date: Wed Apr 28 2004 - 23:58:59 EDT
------------------------------------------------------------
Date: Wed Apr 28 23:58:58 2004, Shift: 16:00-24:00
Supervisor: JHL, with: 
Log book#15 pages 115 - 115

Shift Summary:
------------------------------------------------------------
1. Beam for 2.5 hours with ZDC~280-220/sec
   (Since 21:30-)

2. Data (from run11787 -)

   FS 4deg. 1/1A
   MRS 90 deg. 2B

   Scale down factors have been optimized to miximize trigger2 rate
   to keep DAQ live at ~70-80%. (8.5K Trigger2/hours at this setting)  

3. No new detector/DAQ issues

------------------------------------------------------------

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Wed Apr 28 23:59:14 2004

This archive was generated by hypermail 2.1.8 : Wed Apr 28 2004 - 23:59:35 EDT